How many grams of water can be made when 17.5 moles of HNO3 are consumed


1
Expert's answer
2022-05-04T02:21:03-0400

S + 6HNO3 = H2SO4 + 6NO2 + 2H2O

m(H2O) = (m(HNO3) * 2Mr(H2O)) / 6Mr(HNO3)


m(HNO3) = n(HNO3) * Mr(HNO3)

m(HNO3) = 17.5 mol * 63.01 g/mol = 1102.68 g


m(H2O) = (1102.68 g * (2*18 g/mol)) / 6 * 63.01 g/mol = 105 g